# FairJob: A Real-World Dataset for Fairness in Online Systems

### v1

@@ 1 Introduction
To foster research in fairness within real-world scenarios, we release a large-scale fairness-aware dataset for advertising. The dataset contains pseudononymized users’ context and publisher features that were collected from a job targeting campaign ran for 5 months. The data has been sub-sampled non-uniformly to avoid disclosing business metrics. Feature names have been anonymized for business confidentiality, and their values randomly projected to preserve predictive power while making the recovery of the original features or user context (i.e. re-indentification) practically impossible, with accordance to the privacy-safety measures333’Pseudonymisation’ of data – defined in Article 4(5) of General Data Protection Regulation (GDPR) – means replacing any information which could be used to identify an individual with a pseudonym, or, in other words, a value which does not allow the individual to be directly identified. Re-indentification of the pseudomymized data is impossible due to additional randomization techniques during data anonymization. The original data does not contain any ”special categories” of personal data listed under Article 9 of the GDPR, processing of which is prohibited, except in limited circumstances set out in Article 9 of the GDPR. The publication of the dataset was approved by a Data Protector Officer and Legal professionals.. Although our dataset does not contain explicit sensitive attributes such as gender, it includes a gender proxy derived from non-protected relevant attributes, which we discuss in detail further.
@@ 3.1 Fairness definition
We base our discussion on a counterfactual fairness framework that explains the underlying connections between the variables in the system [Kusner et al., 2017]. Let $A$ denote a protected attribute (can be a set of protected attributes) of an individual, $X$ denote the other observable attributes of any particular individual, $Y$ denote the outcome to be predicted, and let $\hat{Y}$ be a predictor. The predictor takes into account the available data from logs of user interactions with the system and product descriptions and estimates the probability of a positive outcome, i.e. click of the user on the product. The system takes into account the prediction and then shows the best product to the user, which results into possible positive outcome. In our analysis, we are interested in understanding how $A$ and $X$ influence $Y$ and how well our predictor $\hat{Y}$ captures these relationships. Our goal is not just to predict outcomes accurately but also to ensure fairness and mitigate biases in the predictions with respect to $A$. There random variables have the causal relationships that are modelled on Fig. 2 and we discuss further in detail.
@@
Figure 2: Causal graph depicting effects of variables appearing during model training under different constraints.
@@ 3.2 Selection bias in campaign choosing
Note that the recommendation engines predict $\mathbb{P}(Y=1\mid A=a,C=c)$ for a product in the campaign $c$. Thus, if we use prediction bias mitigation techniques while considering data at the campaign level, in the best case scenario, we obtain fair predictions while being unfair outside of campaing, $\mathbb{P}(\hat{Y}=1\mid A=0,C=c)=\mathbb{P}(\hat{Y}=1\mid A=1,C=c)$ and $\mathbb{P}(\hat{Y}=1\mid A=0)\not=\mathbb{P}(\hat{Y}=1\mid A=1)$ Thus, we have to take into account the selection bias to ensure demographic parity introduced in Eq. (2):
$\mathbb{P}(\hat{Y}=1\mid A=1,C=c)\frac{\mathbb{P}(A=1)}{\mathbb{P}(A=1\mid C=c)}=\mathbb{P}(\hat{Y}=1\mid A=0,C=c)\frac{\mathbb{P}(A=0)}{\mathbb{P}(A=0\mid C=c)}.$ (3)
@@ 3.4 Recommendation bias
$U(f)=\mathbb{E}_{I}\mathbb{E}_{X,D\mid I}\bigl{[}\mathbb{I}\left(Y_{D}=1\right)\,\text{rank}_{I}f(X_{I},D)\bigr{]},$ (4)
where $\mathbb{I}(Y_{D}=1)$ is the identity function of a positive outcome (e.g. click) for display $D$. The function $\text{rank}_{D}$ computes the ascending order rank within the set of displays for a banner of impression $I$. This metrics is based on estimation of the positive outcome based on the passed events for chosen users.
@@
$\tilde{U}(f)=\mathbb{E}_{D}\mathbb{E}_{I\mid D}\left[\mathbb{I}\left(Y_{D}=1\right)\,\frac{\mathbb{P}(A=a_{X_{I}})}{\mathbb{P}(A=a_{X_{I}}|C=c)}\,\text{rank}_{I}f(X_{I},D)\right],$ (5)
where $a_{X}$ stands for a gender of a given user $X$. Intuitively, if the prediction is biased with respect to protected attribute $A$, the final prediction $\mathbb{P}(Y=1|A=a)$ is even more biased due to selection bias of with respect to the protected attribute of choosing a campaign $C=c$: $\mathbb{P}(C=c\mid A=a)$. In this case, the prediction model amplifies the existing historical bias. However, we can remove the selection bias by adding weights that correspond to the presence of the protected attribute in the whole population and given the campaign. If the user with protected attribute $A=a$ has lower probability of click, and this group was underrepresented in the campaign $C=c$, i.e. $\mathbb{P}(A=a)>\mathbb{P}(A=a|C=c)$, then in the utility function, the model’s prediction will be higher, by addressing the possible bias due to under-representation in the data. This is our suggested metric to evaluate the recommendation system when the selection bias is present and known such as in the FairJob dataset.
@@ 4 FairJobs dataset
We introduce FairJobs [^7] https://huggingface.co/datasets/criteo/FairJob dataset that contains fairness-aware data from a real-world scenario of advertising. This dataset is intended to learn click predictions models and evaluate by how much their predictions are biased between different gender groups. The dataset consists of 1,072,226 rows that were collected during 5 months of a targeted job campaign [^8] We leave the details on the data collection, feature engineering and privacy-preserving steps in the supplemental material., each row represents a job ad and user features: 20 categorical and 39 numerical features; label click (binary, if the ad was clicked), protected_attribute (binary, proxy for user gender, see below for more thorough explanation), senior (binary, if the job offer was for a senior position), [user_id, impression_id, product_id] are unqiue identifiers of user, impression and product (job ad). More details and dataset statistics are referred to Appendix.
@@
We remark that the proposed gender proxy does not give a definition of the gender. Since we do not have access to the sensitive information, this is the best solution we have identified at this stage to idenitify bias on pseudonymised data, and we encourage any discussion on better approximations. This proxy is reported as binary for simplicity yet we acknowledge gender is not necessarily binary. Although our research focuses on gender, this should not diminish the importance of investigating other types of algorithmic discrimination. While this dataset provides important application of fairness-aware algorithms in a high-risk domain, there are several fundamental limitation that can not be addressed easily through data collection or curation processes. These limitations include historical bias that affect a positive outcome for a given user, as well as the impossibility to verify how close the gender-proxy is to the real gender value. Additionally, there might be bias due to the market unfairness that we explained in Section 3.3. Such limitations and possible ethical concerns about the task should be taken into account while drawing conclusions from the research using this dataset. Readers should not interpret summary statistics of this dataset as ground truth but rather as characteristics of the dataset only.
@@ 5 Empirical observations
Results.
We choose two baseline methods: (i) unfair, that uses all attributes for training, including the protected one; and (ii) unaware, that corresponds to fairness through unawareness, i.e. using all attributes during training except the protected one. We compare the results of these algorithms with a (iii) fair model that is trained without protected attribute with additional "fairness" penalty [Kamishima et al., 2011, Bechavod and Ligett, 2017]. The three models correspond to the situations described in Figure 2. We refer all the reproducibility details and additional experiments to the supplemental material. The resulted prediction can be visually compared in Figure 5.
@@ Results.
Additionally, we propose to restraint an access to the protected attribute and study the trade-off when we train the model on the whole train set but add fairness penalty only for some percentage of train set. In some scenarios, we see improvements in fairness without sacrificing the overall performance. The loss in accuracy due to the imposed fairness constraints is often small as also noted in other works [Celis et al., 2019b]. We explore bias correction techniques tailored to address data limitations and preserve utility in large-scale. We demonstrate how prioritizing fairness in AI not only benefits users by fostering inclusivity but also contributes to the long-term success and ethical integrity of companies. Details and results of the experiments are reported in the supplemental material.
@@ 6 Conclusion
Addressing bias in AI goes beyond mere compliance with legal frameworks like the AI Act; it necessitates proactive measures to detect, prevent, and mitigate biases. Drawing from real-world challenges faced by industries, we highlight the limitations of existing bias mitigation strategies, particularly in environments where access to sensitive user attributes is restricted. We encourage other authors and practitioners to experiment with different AI or Fair AI algorithms on this dataset. We expect that with this work, the quality of evaluation of novel AI methods increases, potentiating the development of the area. Additionally, we hope it encourages other similar relevant datasets to be published from other authors and institutions.
Appendix

### v2

@@ Abstract
The dataset is hosted at https://huggingface.co/datasets/criteo/FairJob. Source code for the experiments is hosted at https://github.com/criteo-research/FairJob-dataset/.
@@ 1 Introduction
To foster research in fairness within real-world scenarios, we release a large-scale fairness-aware dataset for advertising. The dataset contains pseudononymized users’ context and publisher features that were collected from a job targeting campaign ran for 5 months. The data has been sub-sampled non-uniformly to avoid disclosing business metrics. Feature names have been anonymized for business confidentiality, and their values randomly projected to preserve predictive power while making the recovery of the original features or user context (i.e. re-indentification) practically impossible, with accordance to the privacy-safety measures333’Pseudonymisation’ of data – defined in Article 4(5) of General Data Protection Regulation (GDPR) – means replacing any information which could be used to identify an individual with a pseudonym, or, in other words, a value which does not allow the individual to be directly identified. The orginal data is still used in the AdTech company, however, re-indentification of the pseudomymized data is impossible due to additional randomization techniques during data anonymization. The original data does not contain any ”special categories” of personal data listed under Article 9 of the GDPR, processing of which is prohibited, except in limited circumstances set out in Article 9 of the GDPR. The publication of the dataset was approved by a Data Protector Officer and Legal professionals.. Although our dataset does not contain explicit sensitive attributes such as gender, it includes a gender proxy derived from non-protected relevant attributes, which we discuss in detail further.
@@ 3.1 Fairness definition
We base our discussion on a counterfactual fairness framework that explains the underlying connections between the variables in the system [Kusner et al., 2017]. Let $A$ denote a protected attribute (can be a set of protected attributes) of an individual, $X$ denote the other observable attributes of any particular individual, $Y$ denote the outcome to be predicted, and let $\hat{Y}$ be a predictor. The predictor takes into account the available data from logs of user interactions with the system and product descriptions and estimates the probability of a positive outcome, i.e. click of the user on the product. The system takes into account the prediction and then shows the best product to the user, which results into a possible positive outcome. In our analysis, we are interested in understanding how $A$ and $X$ influence $Y$ and how well our predictor $\hat{Y}$ captures these relationships. Our goal is not just to predict outcomes accurately but also to ensure fairness and mitigate biases in the predictions with respect to $A$. These random variables exhibit causal relationships, as modeled in Fig. 2, which we further explore in detail below.
@@
Figure 2: Causal graph depicting effects of variables appearing during model training under different constraints. The arrow between the nodes corresponds to the causal effect. The dashed arrow between $\hat{Y}$ and $X$ can be interpreted as $\hat{Y}$ depends on $X$, but conditionally on $X$, $\hat{Y}$ is independent to its ancestors.
@@ 3.2 Selection bias in campaign choosing
Note that the recommendation engines predict $\mathbb{P}(Y=1\mid A=a,C=c)$ for a product in the campaign $c$. Thus, if we use prediction bias mitigation techniques while considering data at the campaign level, in the best case scenario, we obtain fair predictions while being unfair outside of campaing, $\mathbb{P}(\hat{Y}=1\mid A=0,C=c)=\mathbb{P}(\hat{Y}=1\mid A=1,C=c)$ and $\mathbb{P}(\hat{Y}=1\mid A=0)\not=\mathbb{P}(\hat{Y}=1\mid A=1)$ Thus, we have to take into account the selection bias to ensure demographic parity introduced in Eq. (2). Details on the derivation of the campaign selection bias and its correction are referred to supplemental material.
@@ 3.4 Recommendation bias
$U(f)=\mathbb{E}_{J}\mathbb{E}_{X,D\mid J}\bigl{[}\mathbb{I}\left(Y_{D}=1\right)\,\text{rank}_{J}f(X_{J},D)\bigr{]},$ (3)
where $\mathbb{I}(Y_{D}=1)$ is the identity function of a positive outcome (e.g. click) for display $D$. The function $\text{rank}_{D}$ computes the ascending order rank within the set of displays for a banner $J$. This metrics is based on estimation of the positive outcome based on the passed events for chosen users.
@@
$\tilde{U}(f)=\mathbb{E}_{D}\mathbb{E}_{J\mid D}\left[\mathbb{I}\left(Y_{D}=1\right)\,\frac{\mathbb{P}(A=a_{X_{J}})}{\mathbb{P}(A=a_{X_{J}}|C=c)}\,\text{rank}_{J}f(X_{J},D)\right],$ (4)
where $a_{X}$ stands for a gender of a given user $X$. Intuitively, if the prediction is biased with respect to the protected attribute $A$, the final prediction $\mathbb{P}(Y=1|A=a)$ is even more biased due to selection bias with respect to the protected attribute of choosing a campaign $C=c$: $\mathbb{P}(C=c\mid A=a)$. In this case, the prediction model amplifies the existing historical bias. However, we can remove the selection bias by adding weights that correspond to the presence of the protected attribute in the whole population and given the campaign. If the user with protected attribute $A=a$ has lower probability of click, and this group was underrepresented in the campaign $C=c$, i.e. $\mathbb{P}(A=a)>\mathbb{P}(A=a|C=c)$, then in the utility function, the model’s prediction will be higher, by addressing the possible bias due to under-representation in the data. This is our suggested metric to evaluate the recommendation system when the selection bias is present and known such as in the FairJob dataset.
@@ 4 FairJobs dataset
We introduce FairJobs [^7] https://huggingface.co/datasets/criteo/FairJob dataset that contains fairness-aware data from a real-world scenario of advertising. The intended use of this dataset is to learn click predictions models and evaluate by how much their predictions are biased between different gender groups. The dataset consists of 1,072,226 rows that were collected during 5 months of a targeted job campaign [^8] We leave the details on the data collection, feature engineering and privacy-preserving steps in the supplemental material., each row represents a job ad and user features: 20 categorical and 39 numerical features; label click (binary, if the ad was clicked), protected_attribute (binary, proxy for user gender, see below for more thorough explanation), senior (binary, if the job offer was for a senior position), [user_id, impression_id, product_id] are unqiue identifiers of user, impression and product (job ad). More details and dataset statistics are referred to Appendix.
@@
We remark that the proposed gender proxy does not give a definition of the gender. Since we do not have access to the sensitive information, this is the best solution we have identified at this stage to idenitify bias on pseudonymised data, and we encourage any discussion on better approximations. This proxy is reported as binary for simplicity yet we acknowledge gender is not necessarily binary. Although our research focuses on gender, this should not diminish the importance of investigating other types of algorithmic discrimination. While this dataset provides important application of fairness-aware algorithms in a high-risk domain, there are several fundamental limitation that can not be addressed easily through data collection or curation processes. These limitations include historical bias that affect a positive outcome for a given user, as well as the impossibility to verify how close the gender-proxy is to the real gender value. Additionally, there might be bias due to the market unfairness that we explained in Section 3.3. Such limitations and possible ethical concerns about the task should be taken into account while drawing conclusions from the research using this dataset. Readers should not interpret summary statistics of this dataset as ground truth but rather as characteristics of the dataset only. Additional limitation comes from identifying the senior position label, as the definition of what constitutes a "senior" position can be subjective. We acknowledge that this method may introduce some noise, particularly if job titles are unconventional or if errors occur in categorization. Finally, we remark that in the dataset we assume that each user_id represents a single user; however, we acknowledge that multiple users could share one device, potentially affecting the user features. Additionally, user_id’s are based on the company’s identification technology, which means that a single user could have multiple user_id’s across different browsing sessions. This is one of the complexities inherent in real-world data, particularly in online advertising. Such biases can influence model training, bias evaluation, and bias mitigation efforts.
@@ 5 Empirical observations
Baselines.
We choose two baseline regimes: (i) unfair, that uses all attributes for training, including the protected one; and (ii) unaware, that corresponds to fairness through unawareness, i.e. using all attributes during training except the protected one. We train (i) a Dummy classifier in the unaware regime to obtain the first baseline and (ii) XGB in two regimes (unaware and unfair) to achieve a more reasonable baseline performance, see Table 1. We notice that the Dummy classifier is perfectly fair with respect to demographic parity DP which is reasonable since it did not learn the dependence between features at the label at all, as can be seen from the negative log-likelihood NLLH results. However, the utility metrics $U$ and $\tilde{U}$ of Dummy do not differ much from XGB which is due to very strong imbalance in the data. We remark that $\tilde{U}$ is expectedly higher for fairer (unaware) models, while $U$ is better for the unfair model. There is a slight difference in terms of NLLH and AUC-ROC for XGB models, and higher $U$ corresponds to higher AUC-ROC. We refer to Appendix B and C for more details.
NLLH $\downarrow$ AUC $\uparrow$ DP $\downarrow$ $U$uparrow $\tilde{U}$uparrow
Dummy unaware 0.69239 0.50000 0.00000 0.01009 0.01245
XGB unaware 0.05491 0.75787 0.00278 0.01017 0.01276
XGB unfair 0.05736 0.76201 0.00323 0.01037 0.01236
Fair regime.
To study a possible trade-off between utility and fairness, we use in-processing fairness methods of adding a fairness-inducing penalty to a loss function $\mathcal{L}(\hat{Y},Y)$ during training [Kamishima et al., 2011]:
$\hat{Y}=\arg\min\ \ \mathcal{L}(\hat{Y},Y)+\lambda\cdot\text{Penalty}(\hat{Y},Y,A),$ (5)
where parameter $\lambda$, or fairness_multiplier, controls the trade-off between the model’s predictive accuracy $\mathcal{L}(\hat{Y},Y)$ and fairness. Adjustments on $\lambda$ allows to control the importance of fairness relative to accuracy. These methods remove the influence of protected attribute on the model’s output without restrictions on the data [Kamishima et al., 2011, Bechavod and Ligett, 2017, Mary et al., 2019]. We implement a penalty based on the approach described in Bechavod and Ligett [2017].
We train a logistic regression in the two baseline regimes and compare the results of these algorithms with a (iii) fair model that is trained without protected attribute with an additional fairness-inducing penalty [Kamishima et al., 2011, Bechavod and Ligett, 2017]. The three models correspond to the situations described in Figure 2. We refer all the reproducibility details and additional experiments to Appendix B and C. The resulted prediction can be visually compared in Figure 5.
@@
Fairness-utility trade-off.
Additionally, we propose to restraint an access to the protected attribute and study the trade-off when we train the model on the whole train set but add fairness penalty only for some percentage of train set. In some scenarios, we could see improvements in fairness without sacrificing the overall performance. The loss in accuracy due to the imposed fairness constraints is often small as also noted in other works [Celis et al., 2019b]. We explore bias correction techniques tailored to address data limitations and preserve utility in large-scale. We demonstrate how prioritizing fairness in AI not only benefits users by fostering inclusivity but also contributes to the long-term success and ethical integrity of companies.
@@ 6 Conclusion
Addressing bias in AI goes beyond mere compliance with legal frameworks like the AI Act; it necessitates proactive measures to detect, prevent, and mitigate biases. Drawing from real-world challenges faced by industries, we highlight the limitations of existing bias mitigation strategies, particularly in environments where access to sensitive user attributes is restricted. We encourage other authors and practitioners to experiment with different AI or Fair AI algorithms on this dataset. We argue that specific problems can often be generalized to broader contexts. For example, if our dataset helps identify a method that effectively balances fairness and utility, this method could potentially be applicable to other recommendation systems across various domains. We expect that with this work, the quality of evaluation of novel AI methods increases, potentiating the development of the area, see more details in the broader impact section in Appendix E. Additionally, we hope it encourages other similar relevant datasets to be published from other authors and institutions.
7 Acknowledgements
Federico Pavone has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034255. We also thank Martin Bompaire, David Rhode and Andre Cunha for helpful discussions.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] Section 4, paragraph "Limitations and interpretation".
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [Yes]
    2. (b)
      Did you include complete proofs of all theoretical results? [Yes]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Section Reproducilibity in supplemental meterial
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Section Reproducilibity in supplemental meterial
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] We repeated experiments 10 times with random seeds.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Section Reproducilibity in supplemental meterial
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
8 Checklist for the dataset track
  1. 1.
    Submission introducing new datasets must include the following in the supplementary materials:
    1. (a)
      Dataset documentation and intended uses. Recommended documentation frameworks include datasheets for datasets, dataset nutrition labels, data statements for NLP, and accountability frameworks. A: Available on the dataset page and in the text in paragraph "License and intended use".
    2. (b)
      URL to website/platform where the dataset/benchmark can be viewed and downloaded by the reviewers. A: The dataset is hosted on HuggingFace dataset API: FairJob, the other information can be found in supplemental material.
    3. (c)
      URL to Croissant metadata record documenting the dataset/benchmark available for viewing and downloading by the reviewers. You can create your Croissant metadata using e.g. the Python library available here: https://github.com/mlcommons/croissant The dataset viewer automatically generates the metadata in Croissant format (JSON-LD) for every dataset on the Hugging Face Hub. It lists the dataset’s name, description, URL, and the distribution of the dataset as Parquet files, including the columns’ metadata. The Croissant metadata is available for all the datasets that can be converted to Parquet format.
    4. (d)
      Author statement that they bear all responsibility in case of violation of rights, etc., and confirmation of the data license. A: the statement is in the supplemental material.
    5. (e)
      Hosting, licensing, and maintenance plan. The choice of hosting platform is yours, as long as you ensure access to the data (possibly through a curated interface) and will provide the necessary maintenance. A: all the details are in supplemental material and on the Hugging Face dataset webpage.
  2. 2.
    To ensure accessibility, the supplementary materials for datasets must include the following:
    1. (a)
      Links to access the dataset and its metadata. This can be hidden upon submission if the dataset is not yet publicly available but must be added in the camera-ready version. In select cases, e.g when the data can only be released at a later date, this can be added afterward. Simulation environments should link to (open source) code repositories.
    2. (b)
      The dataset itself should ideally use an open and widely used data format. Provide a detailed explanation on how the dataset can be read. For simulation environments, use existing frameworks or explain how they can be used.
    3. (c)
      Long-term preservation: It must be clear that the dataset will be available for a long time, either by uploading to a data repository or by explaining how the authors themselves will ensure this.
    4. (d)
      Explicit license: Authors must choose a license, ideally a CC license for datasets, or an open source license for code (e.g. RL environments).
    5. (e)
      Add structured metadata to a dataset’s meta-data page using Web standards (like schema.org and DCAT): This allows it to be discovered and organized by anyone. If you use an existing data repository, this is often done automatically.
    6. (f)
      Highly recommended: a persistent dereferenceable identifier (e.g. a DOI minted by a data repository or a prefix on identifiers.org) for datasets, or a code repository (e.g. GitHub, GitLab,…) for code. If this is not possible or useful, please explain why.
    A: All the dataset details including links, metadata, preservation, license, reproducility details are avaiable in the supplemental material. Licence and intended use information is also stated in the main paper text.
  3. 3.
    For benchmarks, the supplementary materials must ensure that all results are easily reproducible. Where possible, use a reproducibility framework such as the ML reproducibility checklist, or otherwise guarantee that all results can be easily reproduced, i.e. all necessary datasets, code, and evaluation procedures must be accessible and documented. A: All details for the reproducibility are available in the supplemental material.
  4. 4.
    For papers introducing best practices in creating or curating datasets and benchmarks, the above supplementary materials are not required.
Supplementary materials